x - y = 3
what does the graph look like

Found 3 solutions by ikleyn, mananth, josgarithmetic:
Answer by ikleyn(53762)

.

Go to web-site

www.desmos.com/calculator

There is free of charge plotting tool there for common use.

Print your equation there.

You will get the plot in the next instance, right in front of your eyes.



Answer by mananth(16949)

you can use any graphing calculator
OR
Plot the points and draw the line
x-y=3
y=x-3
Get y values from different values of x
For
x=1 , y= -2 (1,-2)
x=3 , y=0 (3,0)
x=-3, y=-6 (-3,-6)
Plot the points and draw the line
.





Answer by josgarithmetic(39797)

What does the graph look like?
Find the axis intercept points, plot them, and draw the line!